a moving body A of mass 5kg collides with a stationary body B. If both the bodies move togeather with a speed of 1/3 that of the body A after collision, mass of B is
1
Expert's answer
2012-09-13T11:01:52-0400
Let Ia be the momentum of a moving body A before collision, Iab be the momentum of A and B after collision, Va the velocity of A before collision, Vab velocity of A and B after collision, Ma and Mb the masses of A and B respectively. Then

Ia = Ma*Va
Iab = (Ma+Mb)*Vab
Vab = Va/3
Ia = Iab
Ma*Va = (Ma+Mb)*Vab
Ma*Va = (Ma+Mb)*Va/3
Ma = (Ma+Mb)/3
Ma = Ma/3 + Mb/3
3*Ma = Ma + Mb
2*Ma = Mb
Mb = 2*5 = 10 kg

So, mass of B is 10 kg.